A car is traveling on a road that is perpendicular to a railroad track. When the car is30 meters from the crossing, the car’s new collision detector warns the driver that there is a train 50 meters from the car and heading toward the same crossing. How far is the train from the crossing?
Answer by richwmiller(17219)

a^2+b^2=c^2
30^2+b^2=50^2
900+b^2=2500
b^2=1600
b=40